# Representative Ranking for Deliberation in the Public Sphere

### v1

@@ 10 $12$ $1$ $10$ $2$ $1$ $1$
Figure E.2: Representation when selecting comments using diverse approval with and without a JR constraint (missing feeds).
@@ E.6 Results Per Political Ideology
Figure E.3: Representation of political groups when ranking with the diverse approval $f_{\text{eng}}$
Figure E.4: Representation of political groups when ranking with the diverse approval $f_{\text{DA}}$
Figure E.5: Representation of political groups when ranking with the Perspective API $f_{C}$
@@ E.7 Individual Attributes from Perspective API
Figure E.6: Representation of users when ranking with individual Perspective API attributes, both with the JR constraint ($S^{\text{Greedy}}_{JR}$) and without the JR constraint ($S^{*}$).

### v2

@@ 10 $12$ $1$ $10$ $2$ $1$ $1$
Fig. E.2: Representation when selecting comments using diverse approval with and without a JR constraint (missing feeds).
In Figure E.3, we plot the average score per comment for the set that maximizes the scoring function $f$ as a function of the percentage of users represented by the selected set. This is shown for both the unconstrained optimal set $S^{*}(f)$ and the set $S^{\text{Greedy}}_{JR}(f)$ that satisfies JR and approximately maximizes the score function $f.$ We effectively calculate $\frac{f(S^{*}(f))}{k}$ and $\frac{f(S^{\text{Greedy}}_{JR}(f))}{k}$ respectively. [^19] We acknowledge an anonymous reviewer for suggesting this plot.
Fig. E.3: The score-optimal sets $S^{*}$ and the JR sets $S^{\text{Greedy}}_{JR}$ for each Remesh question (see Table E.1) and score function (engagement, diverse approval, and the Perspective API bridging score). Each set is plotted based on the average score of the items in the set and the percentage of users represented in the set. Compared to the score-optimal sets $S^{*}$, the JR sets $S^{\text{Greedy}}_{JR}$ notably increase the percentage of represented users without significantly reducing the score of the selected items.
@@ E.6 Results Per Political Ideology
Fig. E.4: Representation of political groups when ranking with the diverse approval $f_{\text{eng}}$
Fig. E.5: Representation of political groups when ranking with the diverse approval $f_{\text{DA}}$
Fig. E.6: Representation of political groups when ranking with the Perspective API $f_{C}$
@@ E.7 Individual Attributes from Perspective API
Fig. E.7: Representation of users when ranking with individual Perspective API attributes, both with the JR constraint ($S^{\text{Greedy}}_{JR}$) and without the JR constraint ($S^{*}$).